Find the (a) mean, (b) median, (c) mode, and (d) midrange for the data and then (e) answer the given question. Listed below are foot lengths in inches of randomly selected women in a study of a countrys military in 1988. Are the statistics representative of the current population of all women in that countrys military? 8.6 9.4 8.7 9.4 9.5 9.7 9.0 9.3 9.1 8.7 8.7 a. Find the mean. The mean is □ inches). (Type an integer or a decimal rounded to two decimal places as needed.) b. Find the median. The median isinch(es). (Type an integer or a decimal. Do not round.) c. Find the mode. Select the correct choice below and, if necessary, fill in the answer box to complete your choice O A. The mode(s) is(are) Type an integer or a decimal. Do not round. Use a comma to separate answers as needed.) There is no mode. inch(es). 0 B. d. Find the midrange The midrange is inch(es). Type an integer or a decimal. Do not round.) e. Are the statistics representative of the current population of all women in that countrys military? Choose the best answer below. O A. 0 B. O C. 0 D. Since the sample does not include men, the sample should not be considered to be representative of the population. Since the sample is random and the sample size is greater than 10, the sample can be considered to be representative of the population. Since the sample is not a random sample, it should not be considered to be representative of the population. Since the measurements were made in 1988, they are not necessarily representative of the current population of all women in the countrys military Click to select your answer(s)
Answer #1

ANSWER:-

Given data,

8.6 9.4 8.7 9.4 9.5 9.7 9.0 9.3 9.1 8.7 8.7

a).Here, we have to find the mean:-

Mean=(8.6+9.4+8.7+9.4+9.5+9.7+9.0+9.3+9.1+8.7+8.7)/11

=100.1/11

=9.1

The mean is 9.1 inches

b).Here, we have to find the median:-

First we arrange the given values in ascending order then,

8.6,8.7,8.7,8.7,9.0,9.1,9.3,9.4,9.4,9.5,9.7

The median is middle term of the ascending values

The median is 9.1 inches

c).Here, we have to calculate the mode:-

8.6,8.7,8.7,8.7,9.0,9.1,9.3,9.4,9.4,9.5,9.7

more times repeated number

The mode is 8.7 inches.

d).Here, we have to find the mid range:-

Range=largest value -Smallest value

=9.7-8.6

=1.1

The mid range is 1.1 inches.

e).Option d is correct
Answer #2
No thanks. Fuck that noise
source: 18
answered by: Ashley